Open Source Writers Group: Document Licensing Policy

1. Purpose of this Document

This document has been written in order to simplify the process of deciding which licenses are allowable for documents that are to be included as part of the Open Source Writers Group documentation set.

Rather than having a license-related debate every time a new open-source or open-content license is introduced, we here set out the basic "baseline" freedoms that each license must include in order to be "free enough" to be appropriate for use with documents that are associated with the OSWG. This policy statement has been made as simple as possible in order to avoid any future confusion or conflict about what this policy entails. Under no circumstances, however, should this policy be considered as "written in stone". If there is sufficient demand by members of the OSWG, this policy can and will be modified.

In this document, "documentation set" is a general term that applies to the full collection of documentation, papers, articles, books, FAQs, HOWTOs, and other written material that is associated with the Open Source Writers Group. Excepting for unusual circumstances, the full OSWG documentation set will be included within the OSWG CVS system available on the OSWG server. Where possible, OSWG documents will be made available in multiple formats including the original DocBook SGML source.

2. Basic Licensing Requirements

Open Source Writers Group: Document Licensing Policy

Any document that is to be part of the OSWG documentation set must be released under an open-source or open-content license that allows that document to be, at minimum:

1. freely reformatted

2. freely distributed

There are a wide variety of licenses available under which documents can be released, ranging across a whole spectrum of possible levels of "freedom". The two requirements listed above mark a point on that spectrum, and any license that is "more free" (has fewer licensing restrictions) is appropriate for use in the OSWG documentation set. Any license that is "less free" (has greater licensing restrictions) is not appropriate.

3. Further Discussion

To put these licensing requirements in more concrete terms, if you wish to have your document included in the OSWG documentation set, you must release that document under a license which, at very least, allows any other person to take that document, modify its formatting, and redistribute it in any form, including forms that are intended for sale.

So, if a publishing company so desires, they can take the entire contents of the OSWG documentation set, publish those documents as a book, and then sell that book. This is akin to the creation and sale of a Linux distribution by companies such as Red Hat or Caldera – they are able to take Linux and a variety of other open-source programs, package those together as a distribution, and sell that package.

If you wish to release your document under a license that has fewer licensing restrictions, that document will be allowed into the OSWG documentation set. For example, if you wish to allow others to freely modify, update, and translate your document, you can release it under a license that allows for this. This type of license is allowed (and, in fact, strongly encouraged) for documents in the OSWG documentation set.

On the other hand, documents released under licenses that are more restrictive are not allowed to be part of the OSWG documentation set. For example, a license that allows for the free reformatting and redistribution of all electronic versions of a document, but which restricts the publication and distribution of that document in print/hardcopy format, is not an appropriate license for inclusion in the OSWG documentation set. You cannot include documents in the OSWG documentation set over which you wish to reserve exclusive print publication rights. One example of a license that is not allowable for this reason is the Open Publication License (OPL – http://www.opencontent.org/openpub (http://www.opencontent.org/openpub/)) when that license includes Option B.

In the first requirement listed above – that a license must allow for a document to be freely reformatted – "formatting" here refers to the presentation-related or metadata-related aspects of that document. Formatting does not include the textual or graphical content of the document.

4. A list of allowable licenses, incomplete

Open Publication License - http://www.opencontent.org/openpub/

Note: OPL cannot include Option B

Need more allowable licenses and refs here.